OpenStudy (anonymous):

Help with precalc identity type question? Medal and fan If sinx= -3/5 and x is in quadrant 3, then tan2x = A. 24/7 B. 12/7 C. -12/7 D. -24/7 Through calculation, I came up with C. Is that correct?

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

I was thinking 4 quadrant sorry. So it would be 12/7 yeah?

rishavraj (rishavraj):

its the wrong reply

zepdrix (zepdrix):

Hmm what do we want to use here? Tangent Double Angle Identity I guess?

OpenStudy (anonymous):

Yep, 2tanx/ 1-tan^2x

zepdrix (zepdrix):



rishavraj (rishavraj):

@seehearcreate i guess its wrong question..all bcoz if x lies in 3rd quadrant then sin and cos both are supposed to have negative value...but according to question x comes to be positive..its impossible

zepdrix (zepdrix):

There's a triangle in the third quadrant. Understand how to label your sides using Sin x? :o

rishavraj (rishavraj):

its cos x comes to be +ve

OpenStudy (anonymous):

Not quite....

zepdrix (zepdrix):

@rishavraj , the question is asking for a completely different angle. 2x instead of x. I think it's probably not impossible :o Let's see how it works out.

zepdrix (zepdrix):

\[\Large\rm \sin x=\frac{opposite}{hypotenuse}=-\frac{3}{5}\]

zepdrix (zepdrix):

Tip: The hypotenuse should never be the `negative` side. So put the negative with the 3.

OpenStudy (anonymous):

Ohhh, yeah I knew that sorry, I thought you meant something else in regards to the triangle in the unit circle.

zepdrix (zepdrix):

oh :3